# FinCoT: Grounding Chain-of-Thought in Expert Financial Reasoning

### v1

@@ Abstract
This paper presents FinCoT, a structured chain-of-thought (CoT) prompting approach that incorporates insights from domain-specific expert financial reasoning to guide the reasoning traces of large language models. We investigate that there are three main prompting styles in FinNLP: (1) standard prompting—zero-shot prompting; (2) unstructured CoT—CoT prompting without an explicit reasoning structure, such as the use of tags; and (3) structured CoT prompting—CoT prompting with explicit instructions or examples that define structured reasoning steps. Previously, FinNLP has primarily focused on prompt engineering with either standard or unstructured CoT prompting. However, structured CoT prompting has received limited attention in prior work. Furthermore, the design of reasoning structures in structured CoT prompting is often based on heuristics from non-domain experts. In this study, we investigate each prompting approach in FinNLP. We evaluate the three main prompting styles and FinCoT on CFA-style questions spanning ten financial domains. We observe that FinCoT improves performance from 63.2% to 80.5% and Qwen-2.5-7B-Instruct from 69.7% to 74.2%, while reducing generated tokens eight-fold compared to structured CoT prompting. Our findings show that domain-aligned structured prompts not only improve performance and reduce inference costs but also yield more interpretable and expert-aligned reasoning traces.
@@ 1 Introduction
Financial decision–making including such as stochastic modeling, risk assessment, portfolio optimization, and algorithmic trading (Markowitz, 1952; Black and Scholes, 1973a; Rockafellar and Uryasev, 2000; Avellaneda and Stoikov, 2008). These tasks demand precise mathematics and domain-specific reasoning. Recent advances in large foundation models for finance, such as the multimodal FinTral (Bhatia et al., 2024) and language-centric Fin-R1 (Liu et al., 2025), demonstrate progress, but still face challenges in interpretability and domain alignment (Nie et al., 2024; Arya.ai, 2024; Lee et al., 2025). These shortcomings motivate stricter control over a model’s intermediate reasoning path, which we explore via prompt design.
Prompting guides LLM reasoning without extra training. Methods such as Chain-of-Thought (Wei et al., 2023), Code Prompting (Hu et al., 2023), Plan-and-Solve (Wang et al., 2023), and Self-Reflection (Renze and Guven, 2024) encourage stepwise thinking but remain domain-agnostic. In finance this can lead to omissions in valuation checks or confusion between basis points and percentages. Yet real-world analysis follows well-defined workflows-valuation, discounting, portfolio attribution-that depend on explicit intermediate structure. Embedding such structure in the prompt helps the model verify units, formulas, and boundary conditions, improving interpretability and alignment with expert practice.
@@ 2.1 Prompt Engineering
In addition to these, other prompting techniques have emerged, such as Tree of Thoughts (ToT) (Yao et al., 2023), which explores multiple reasoning paths via tree-structured search; Graph of Thoughts (GoT) (Besta et al., 2024), which frames reasoning as a graph with LLM-generated nodes and edges for flexible information flow. These methods enhance expressiveness for general tasks, they are not tailored for finance, which requires mathematical precision and domain-specific constraints. We adopt the template from Appendix Listing 2 CoT Callanan et al. (2024) as the default prompt for this baseline.
@@
ST-CoT augments a prompt with tags such as <thinking> and <output> that break the model’s response into explicit, modular stages, promoting incremental, easily replaceable reasoning blocks. This tag-driven format has already appeared in open-source trials. [^1] https://gist.github.com/davidmezzetti/1fac6bd406857431f1cdc74545bdfba9 Figure 1 visually contrasts ST-CoT with SP, UST-CoT, and FinCoT.
@@ 2.2 LLMs in Domain-Specific Financial Reasoning
Methods adapt models via supervised fine-tuning with curated QA/classification datasets (Ma et al., 2023; Harsha et al., 2025) or continued pretraining on domain-specific corpora (Yang et al., 2020; Lee et al., 2024; Bhatia et al., 2024; Ke et al., 2025; Liu et al., 2025). While effective, these approaches rely on labeled data and general language modeling objectives, lacking explicit structuring of financial reasoning–thus limiting interpretability and alignment with expert logic.
@@ 3 FinCoT: Augmenting CoT with Structured Financial Expertise
We introduce FinCoT (Financial Chain-of-Thought), a structured prompting framework that enhances LLM reasoning in specialized financial domains. Building upon ST-CoT approaches, FinCoT explicitly embeds expert-derived problem-solving methodologies directly into prompts, guiding LLMs to follow domain-specific reasoning pathways without requiring model fine-tuning. Figure 2 illustrates the FinCoT architecture, which integrates expert-guided reasoning layers and reflective validation to improve performance in financial tasks.
@@ 3.1 The FinCoT Prompt Framework
    System: A single, top-level message that frames the task (e.g., “You are a CFA candidate; treat the following as a finance question").
@@
    Expert Reasoning Blueprint (via Mermaid Diagram (Sveidqvist and contributors, 2025): A domain-specific, embedded expert blueprint with Mermaid diagram concerning the generation of diagrams (see §4), serve as a "Hint" within the context of the prompt. This blueprint explicitly outlines the recommended step-by-step problem-solving strategy for the given financial domain and is curated through a systematic process detailed in section 3.2 to ensure consistency and domain alignment.
@@ 3.2 Embedding Expert Blueprints
To construct expert blueprints for each financial domain, we implement a hybrid pipeline combining machine assistance and human judgment, as illustrated in Figure 3. The process includes the following stages:
@@
    Mermaid Diagram Generation: This refined workflow is translated into a Mermaid diagram (Bari et al., 2024) using its text-based syntax. We selected Mermaid due to its LLM prompt compatibility and clear visual guidance aligning with the FinCoT prompt. The diagrams are constructed based on the source content validated and synthesized first in 2, and then applied to each financial domain, with the entire collection organized and described in Appendix A as reference blueprints.
@@ 4 Experimental Setup
We selected the Qwen family of models due to their strong baseline performance in zero-shot financial reasoning. In preliminary evaluations, Qwen2.5-7B-Instruct achieved 69.7% accuracy on standard task prompts, substantially outperforming LLaMA3.1-8B (46.3%), motivating its use as our investigating model family.
We evaluated both Qwen3-8B-Base and Qwen3-8B in two configurations: (i) their default instruction-tuned format (Qwen3-8B) and (ii) a its reasoning mode (Qwen3-8B (Thinker)). This setup enables direct comparison between inherent model capabilities and those induced through structured prompting.
For a more comprehensive comparison, we also include Fin-R1333https://huggingface.co/SUFE-AIFLM-Lab/Fin-R1, an open-source finance-specific model fine-tuned on Qwen2.5-7B-Instruct into a finance reasoning model. This provides a baseline to evaluate the impact of our prompting strategies beyond a simple instruction tuning variant of the same model.
All experiments used a maximum sequence length of 16,384 tokens. Following best practices for decoding stability (Du et al., 2025), we set the generation temperature to 0.2 to encourage focused and consistent outputs under evaluation conditions.
@@
Our study evaluates the effectiveness of FinCoT against three baseline prompting strategies: SP, UST-CoT, and ST-CoT, which were detailed in Section 2. For clarity in this section:
@@
  • FinCoT: Utilizing zero-shot strategy, this method integrates expert domain templates to refine prompts. A Mermaid diagram acts as a “Hint” guiding the model’s structured analysis of financial issues with pertinent domain insights.
While all methods are zero-shot (meaning that no prior solved examples are shown), FinCoT uniquely provides the model with a rich, structured expert-guided thinking steps as a Mermaid diagram to help solve the current question.
@@
To assess financial reasoning ability, this study uses questions that mirror the rigor of the Chartered Financial Analyst (CFA) examinations. Our evaluation benchmark consists of 1,032 multiple-choice questions selected from the larger FinEval collection, with these specific questions originating from the CFA-Easy dataset (Ke et al., 2025). This curated set allows for a clear evaluation of theoretical and practical financial knowledge. For fine-grained analysis, each question in our benchmark was classified into one of the ten standard CFA domains. This domain classification was performed by GPT-4o using a dedicated classification prompt, as detailed in the Appendix B.3. The resulting domain distribution is visualized in  Figure 4 to support the grouping and evaluation of downstream tasks.
@@
The primary performance measure is accuracy, defined as the percentage of questions for which the model’s predicted choice (A, B, or C) matches the ground-truth answer. To evaluate response efficiency, we also report the average output token count per question, which reflects verbosity and computational cost. In addition, we perform significance testing to determine whether accuracy differences between prompting strategies are statistically significant, we employ a paired bootstrap test (Efron and Tibshirani, 1994). Given two sets of binary correctness scores over the same question set, we perform $B=10{,}000$ bootstrap resamples with replacement and compute the distribution of accuracy differences. From this, we report the mean difference ($\Delta$), 95% confidence interval (CI), and two-tailed $p$-value. This method accounts for the paired nature of the evaluation, where each question serves as its own control, providing a robust statistical basis for comparison.
Furthermore, we also measure the proportion of financial domains where a prompting strategy achieves at least a 1% accuracy improvement over the standard prompt, indicating the breadth of positive impact across domain tasks. Finally, we report the average accuracy improvement (in percentage points) across all financial domains, summarizing overall prompting effectiveness.
@@ 5.1 Financial Reasoning Performance
Prompt Accuracy (%) Average Output Tokens (k)
Qwen2.5-7B Instruct Qwen3-8B Base Qwen3-8B Qwen3-8B (Thinker) Fin-R1 Fin-R1 (Strict Prompt) Qwen2.5-7B Instruct Qwen3-8B Base Qwen3-8B Qwen3-8B (Thinker) Fin-R1 Fin-R1 (Strict Prompt)
SP 69.67 63.18 74.42 88.18 65.70 73.16 0.05* 0.89 0.32 1.52 0.88 2.53
UST-CoT 75.68* 72.58 82.36 89.05* 75.19 74.52* 0.28 0.31* 0.46 1.50 0.58* 2.42
ST-CoT 74.52 78.49 81.01 88.18 74.32 72.48 0.22 3.42 0.25* 1.35 22.20 2.55
FinCoT (All Blueprints) 74.22 80.52* 81.10 87.21 75.78* 73.26 0.29 0.38 0.36 1.23* 1.92 2.41*
@@
FinCoT (Economics) 73.26 79.26 79.55 86.92 73.45 72.00 0.38 0.99 0.39 1.25 2.01 2.48
FinCoT (FixedIncome) 73.35 78.88 80.81 87.21 74.22 74.90 0.27 4.55 0.30 1.24 2.31 2.32
FinCoT (Quant.Meth.) 75.19 80.14 80.91 87.79 74.90 72.48 0.27 3.07 0.31 1.22 2.17 2.62
FinCoT (EquityInvest.) 74.22 79.26 80.52 86.72 74.42 72.38 0.31 7.18 0.37 1.19 2.16 2.78
FinCoT (Port.Manage.) 74.13 80.72 80.91 86.92 75.00 72.87 0.26 0.56 0.30 1.20 2.14 2.68
FinCoT (Derivatives) 73.64 79.84 80.81 87.21 76.16 72.29 0.30 0.42 0.39 1.24 2.05 2.39
FinCoT (Fin. Reporting) 73.84 79.07 81.10 87.02 72.87 71.90 0.28 0.93 0.33 1.19 2.13 2.55
FinCoT (Alter.Invest.) 74.90 78.97 79.94 87.50 74.90 73.64 0.26 0.50 0.38 1.23 2.16 2.42
FinCoT (Corp.Issuers) 74.32 79.26 79.36 87.02 75.58 72.00 0.26 1.18 0.32 1.16 2.08 2.64
Table 1: Comparison of prompting techniques: accuracy (%) and average output token length. ‘FinCoT (All Blueprints)’ simultaneously applies expert reasoning blueprints from all CFA domains, while ‘FinCoT (DomainName)’ (e.g., Economics) applies domain-specific blueprints individually. Results for Fin-R1 utilize its strict fine-tuning prompt [^5] We use the model’s strict fine-tuning prompt from https://github.com/SUFE-AIFLM-Lab/Fin-R1. This prompt is aligned with the ST-CoT prompt and is integrated into each of the other prompting techniques.. (*) Indicates that prompts accuracy improvement over the best-performing all for each model is statistically significant ($p<0.05$), based on paired bootstrap testing.
@@
Footnote 5 summarizes the prompting strategies evaluated in this study, focusing on their accuracy and output efficiency. It contrasts standard prompting (SP), UST-CoT, and FinCoT across models, including Thinker variants, and presents a compute scaling analysis that elucidates trade-offs between token usage and performance-highlighting prompt strategies that enable efficient reasoning with reduced inference cost.
FinCoT (All Blueprints) achieves robust and statistically significant accuracy gains over SP across multiple pretrained models-most notably on Qwen3-8B-Base ($\Delta=17.35\,$pp, 95% CI [15.02, 19.77], $p<0.001$) and Qwen2.5-7B-Instruct ($\Delta=4.55\,$pp, CI [3.29, 5.91], $p<0.001$). Even Fin-R1 (Strict Prompt), a domain-specialized model, benefits from blueprint prompting ($\Delta=1.26\,$pp, CI [0.58, 1.94], $p<0.001$), suggesting that structured financial reasoning complements domain-specific fine-tuning and improves generalization.
In contrast, instruction-tuned models such as Qwen3-8B (Thinker) and Fin-R1 (Strict) exhibit smaller or non-significant gains (e.g., SP vs. FinCoT on Fin-R1 (Strict): $\Delta=-0.10\,$pp, $p=0.73$), indicating that instruction tuning may already encode similar reasoning structures, thereby limiting the marginal utility of blueprint-based prompts.
Pretrained LLMs:
FinCoT yields large and consistent improvements on Qwen3-8B-Base in our experiments, underscoring the value of external domain-specific guidance for models without embedded financial expertise.
Instruction-Tuned Variants:
For models like Qwen2.5-7B-Instruct and Qwen3-8B (Thinker), which have undergone instruction tuning, FinCoT provides minimal additional gains-suggesting overlap between internal tuning objectives and blueprint logic.
Fine-Tuned Financial Model:
Even a domain-specialized model such as Fin-R1 benefits from blueprint prompting, confirming that FinCoT complements rather than duplicates the effects of fine-tuning.

### v2

@@ Abstract
This paper presents FinCoT, a structured chain-of-thought (CoT) prompting framework that embeds domain-specific expert financial reasoning blueprints to guide large language models’ behaviors. We identify three main prompting styles in financial NLP (FinNLP): (1) standard prompting (zero-shot), (2) unstructured CoT (free-form reasoning), and (3) structured CoT (with explicitly structured reasoning steps). Prior work has mainly focused on the first two, while structured CoT remains underexplored and lacks domain expertise incorporation. Therefore, we evaluate all three prompting approaches across ten CFA-style financial domains and introduce FinCoT as the first structured finance-specific prompting approach incorporating blueprints from domain experts. FinCoT improves the accuracy of a general-purpose model, Qwen3-8B-Base, from 63.2% to 80.5%, and boosts Fin-R1 (7B), a finance-specific model, from 65.7% to 75.7%, while reducing output length by up to 8.9$\times$ and 1.16$\times$ compared to structured CoT methods, respectively. We find that FinCoT proves most effective for models lacking financial post-training. Our findings show that FinCoT does not only improve performance and reduce inference costs but also yields more interpretable and expert-aligned reasoning traces.
@@ 1 Introduction
Financial decision–making, such as stochastic modeling, risk assessment, portfolio optimization, and algorithmic trading (Markowitz, 1952; Black and Scholes, 1973a; Rockafellar and Uryasev, 2000; Avellaneda and Stoikov, 2008), demands precise mathematics and domain-specific reasoning Lewkowycz et al. (2022); Wen and Zhang (2025). Recent advances in large foundation models for finance, such as the multimodal FinTral (Bhatia et al., 2024) and language-centric Fin-R1 (Liu et al., 2025), demonstrate progress but still face challenges in interpretability and domain alignment (Nie et al., 2024; Arya.ai, 2024; Lee et al., 2025). Accordingly, these shortcomings motivate stricter control over a model’s intermediate reasoning path, which we explore via prompt design.
Prompting guides LLM reasoning without extra training. Methods such as Chain-of-Thought (Wei et al., 2023), Code Prompting (Hu et al., 2023), Plan-and-Solve (Wang et al., 2023), and Self-Reflection (Renze and Guven, 2024) encourage stepwise thinking but remain domain-agnostic. In finance, this can lead to omissions in valuation checks or confusion between basis points and percentages. Yet real-world analysis follows well-defined workflows—valuation, discounting, portfolio attribution—that depend on explicit intermediate structure. Embedding such structure in the prompt helps the model verify units, formulas, and boundary conditions, improving interpretability and alignment with expert practice.
@@ 2.1 Prompt Engineering
In addition to these, other prompting techniques have emerged, such as Tree of Thoughts (ToT) (Yao et al., 2023), which explores multiple reasoning paths via tree-structured search; Graph of Thoughts (GoT) (Besta et al., 2024), which frames reasoning as a graph with LLM-generated nodes and edges for flexible information flow. These methods enhance expressiveness for general tasks; they are not tailored for finance, which requires mathematical precision and domain-specific constraints. We adopt the template from Appendix Listing 2 CoT Callanan et al. (2024) as the default prompt for this baseline.
@@
ST-CoT augments a prompt with tags such as <thinking> and <output> that break the model’s response into explicit, modular stages, promoting incremental, easily replaceable reasoning blocks. This tag-driven format has already appeared in open-source trials. [^1] https://gist.github.com/davidmezzetti/1fac6bd406857431f1cdc74545bdfba9 Figure˜1 visually contrasts ST-CoT with SP, UST-CoT, and FinCoT.
@@ 2.2 LLMs in Domain-Specific Financial Reasoning
Methods adapt models via supervised fine-tuning with curated QA/classification datasets (Ma et al., 2023; Harsha et al., 2025) or continued pretraining on domain-specific corpora (Yang et al., 2020; Lee et al., 2024; Bhatia et al., 2024; Ke et al., 2025; Liu et al., 2025). While effective, these approaches rely on labeled data and general language modeling objectives, lacking structuring of financial reasoning, thus limiting interpretability and alignment with expert logic.
@@ 3 FinCoT: Augmenting CoT with Structured Financial Expertise
We introduce FinCoT (Financial Chain-of-Thought), a structured prompting framework that enhances LLM reasoning in specialized financial domains. Building upon ST-CoT approaches, FinCoT explicitly embeds expert-derived problem-solving methodologies directly into prompts, guiding LLMs to follow domain-specific reasoning pathways without requiring model fine-tuning. Figure˜2 illustrates the FinCoT architecture, which integrates expert-guided reasoning layers and reflective validation to improve performance in financial tasks.
@@ 3.1 The FinCoT Prompt Framework
    System: A single, top-level message that frames the task (e.g., “You are a CFA candidate; treat the following as a finance question”).
@@
    Expert Reasoning Blueprint (via Mermaid Diagram (Sveidqvist and contributors, 2025): A domain-specific, embedded expert blueprint with Mermaid diagram concerning the generation of diagrams (see §4), serve as a "Hint" within the context of the prompt. This blueprint explicitly outlines the recommended step-by-step problem-solving strategy for the given financial domain and is curated through a systematic process detailed in section˜3.2 to ensure consistency and domain alignment.
@@ 3.2 Embedding Expert Blueprints
To construct expert blueprints for each financial domain, we implement a hybrid pipeline combining machine assistance and human judgment, as illustrated in Figure˜3. The process includes the following stages:
@@
    Mermaid Diagram Generation: This refined workflow is translated into a Mermaid diagram (Bari et al., 2024) using its text-based syntax. We selected Mermaid due to its LLM prompt compatibility and clear visual guidance aligning with the FinCoT prompt. The diagrams are constructed based on the source content validated and synthesized first in 2, and then applied to each financial domain, with the entire collection organized and described in Appendix˜A as reference blueprints [^3] The resulting expert blueprints are reviewed for conceptual consistency and practical correctness (but not guaranteed precision) by CFA Level III charterholders..
@@ 4 Experimental Setup
We selected the Qwen family of models due to their strong baseline performance in zero-shot financial reasoning. In preliminary evaluations, Qwen2.5-7B-Instruct achieved 69.7% accuracy on standard task prompts, substantially outperforming Llama3.1-8B-Instruct (46.3%), motivating its use as our primary model family. To evaluate the impact of both instruction tuning and domain-specific adaptation, we compare two model groups. (A) General-purpose foundation models
  • Qwen2.5-7B (pretrained model) vs. Qwen2.5-7B-Instruct: to assess the effect of instruction tuning on a strong general-purpose foundation model.
  • Qwen3-8B-Base (pretrained model) vs. Qwen3-8B, Qwen3-8B (Thinker): to isolate the impact of ST-CoT prompting within the same architecture.
  • Gemma-3-12B-IT (Text-only): an instruction-tuned model recently released, comparable in size to Qwen3-8B. It achieved 52.81% accuracy on the Flare-CFA benchmark, outperforming Llama3.1-8B-Instruct, and serves as a competitive baseline.
(B) Financial-specific reasoning models
  • Fin-R1 (7B): adapted from Qwen2.5-7B-Instruct using supervised and reinforcement learning on a financial dataset distilled from DeepSeek-R1 (Liu et al., 2025).
  • DianJin-R1-7B: fine-tuned from Qwen2.5-7B-Instruct using CFLUE, FinQA, and CCC, with GRPO to improve domain-specific reasoning (Zhu et al., 2025).
  • Fin-o1-8B: built on Qwen3-8B and trained on the FinCoT [^4] The FinCoT dataset is constructed by TheFinAI and publicly available at https://huggingface.co/datasets/TheFinAI/FinCoT. It combines financial QA datasets such as FinQA, ConvFinQA, TATQA, DocMath-Eval, Econ-Logic, BizBench-QA, and DocFinQA, with GPT-4o-generated reasoning traces to enhance structured financial question answering. Note that this dataset is not derived from our prompting approach. corpus using SFT and GRPO, setting a strong benchmark in financial reasoning (Qian et al., 2025).
All experiments used a maximum sequence length of 16.384k tokens. Following best practices for decoding stability (Du et al., 2025), we set the generation temperature to 0.2 to encourage focused and consistent outputs under evaluation conditions.
@@
Our study evaluates the effectiveness of FinCoT against three baseline prompting strategies: SP, UST-CoT, and ST-CoT, which were detailed in Section˜2. For clarity in this section:
@@
  • FinCoT: A zero-shot prompting method that integrates expert domain templates. Each prompt includes a Mermaid diagram as a “Hint” to guide structured financial reasoning via relevant domain insights.
While all methods operate in a zero-shot setting, FinCoT uniquely injects expert-guided structure through diagrams. Recent studies suggest that even large reasoning models may struggle with instruction-following when overloaded with reasoning cues (Li et al., 2025; Fu et al., 2025; Jang et al., 2025; Yao et al., 2025), though this remains underexplored in financial contexts. We thus evaluate whether CoT-style prompts (UST-CoT, ST-CoT, FinCoT) enhance instruction-following compared to SP.
@@
To assess financial reasoning, we use 1.032k multiple-choice questions from the CFA-Easy subset of FinEval (also referred to as Flare-CFA), originally introduced by Ke et al. (2025). This curated set reflects the rigor of CFA exams and enables evaluation across both theoretical and practical domains. Each question is categorized into one of ten CFA domains using GPT-4o with a dedicated classification prompt (see Appendix B.3), and Figure C shows the resulting domain distribution.
@@
We report accuracy as the metric, defined as the percentage of questions where the model’s prediction matches the ground truth. To assess response efficiency, we also report the average output length (in tokens) across questions. For statistical significance, we use a paired bootstrap test (Efron and Tibshirani, 1994) with $B{=}10k$ resamples over binary correctness scores, reporting the mean difference ($\Delta$), 95% confidence interval, and $p$-value. Additionally, we measure the proportion of financial domains where a method improves accuracy by at least 1% over SP and compute the average domain-wise accuracy gain.
@@ 5.1 Financial Reasoning Performance
Prompt Accuracy (%)
General models Financial models
Qwen2.5-7B
Qwen2.5-7B
Instruct
Qwen3-8B
Base
Qwen3-8B
Qwen3-8B
Gemma-3-12B
IT
Fin-R1
7B
DianJin-R1
7B
Fin-o1
8B
SP 54.07 69.67 63.18 74.42 88.18 52.81 65.70 78.39 79.65
UST-CoT 67.83 ($\uparrow$13.76) 75.68* ($\uparrow$6.01) 72.58 ($\uparrow$9.40) 82.36* ($\uparrow$7.94) 89.05* ($\uparrow$0.87) 77.81* ($\uparrow$25.00) 75.19 ($\uparrow$9.49) 67.73 ($\downarrow$10.66) 79.36 ($\downarrow$0.29)
ST-CoT 70.35* ($\uparrow$16.28) 74.52 ($\uparrow$4.85) 78.49 ($\uparrow$15.31) 81.01 ($\uparrow$6.59) 88.18 76.74 ($\uparrow$23.93) 74.32 ($\uparrow$8.62) 68.80 ($\downarrow$9.59) 78.39 ($\downarrow$1.26)
FinCoT 62.02 ($\uparrow$7.95) 74.22 ($\uparrow$4.55) 80.52* ($\uparrow$17.34) 81.10 ($\uparrow$6.68) 87.21 ($\downarrow$0.97) 75.58 ($\uparrow$22.77) 75.78* ($\uparrow$10.08) 79.75* ($\uparrow$1.36) 77.23 ($\downarrow$2.42)
@@
Economics 69.09 ($\uparrow$15.02) 73.26 ($\uparrow$3.59) 79.26 ($\uparrow$16.08) 79.55 ($\uparrow$5.13) 86.92 ($\downarrow$1.26) 74.61 ($\uparrow$21.80) 73.45 ($\uparrow$7.75) 55.52 ($\downarrow$22.87) 78.00 ($\downarrow$1.65)
FixedIncome 68.12 ($\uparrow$14.05) 73.35 ($\uparrow$3.68) 78.88 ($\uparrow$15.70) 80.81 ($\uparrow$6.39) 87.21 ($\downarrow$0.97) 76.45 ($\uparrow$23.64) 74.22 ($\uparrow$8.52) 66.86 ($\downarrow$11.53) 76.74 ($\downarrow$2.91)
Quant.Meth. 68.02 ($\uparrow$13.95) 75.19 ($\uparrow$5.52) 80.14 ($\uparrow$16.96) 80.91 ($\uparrow$6.49) 87.79 ($\downarrow$0.39) 75.68 ($\uparrow$22.87) 74.90 ($\uparrow$9.20) 65.79 ($\downarrow$12.60) 77.42 ($\downarrow$2.23)
EquityInvest. 69.09 ($\uparrow$15.02) 74.22 ($\uparrow$4.55) 79.26 ($\uparrow$16.08) 80.52 ($\uparrow$6.10) 86.72 ($\downarrow$1.46) 76.45 ($\uparrow$23.64) 74.42 ($\uparrow$8.72) 62.31 ($\downarrow$16.08) 78.68 ($\downarrow$0.97)
Port.Manage. 67.54 ($\uparrow$13.47) 74.13 ($\uparrow$4.46) 80.72 ($\uparrow$17.54) 80.91 ($\uparrow$6.49) 86.92 ($\downarrow$1.26) 77.03 ($\uparrow$24.22) 75.00 ($\uparrow$9.30) 62.02 ($\downarrow$16.37) 76.55 ($\downarrow$3.10)
Derivatives 68.90 ($\uparrow$14.83) 73.64 ($\uparrow$3.97) 79.84 ($\uparrow$16.66) 80.81 ($\uparrow$6.39) 87.21 ($\downarrow$0.97) 77.23 ($\uparrow$24.42) 76.16 ($\uparrow$10.46) 71.80 ($\downarrow$6.59) 79.94 ($\uparrow$0.29)
Fin. Reporting 69.28 ($\uparrow$15.21) 73.84 ($\uparrow$4.17) 79.07 ($\uparrow$15.89) 81.10 ($\uparrow$6.68) 87.02 ($\downarrow$1.16) 75.87 ($\uparrow$23.06) 72.87 ($\uparrow$7.17) 62.69 ($\downarrow$15.70) 77.23 ($\downarrow$2.42)
Alter.Invest. 68.99 ($\uparrow$14.92) 74.90 ($\uparrow$5.23) 78.97 ($\uparrow$15.79) 79.94 ($\uparrow$5.52) 87.50 ($\downarrow$0.68) 76.94 ($\uparrow$24.13) 74.90 ($\uparrow$9.20) 56.98 ($\downarrow$21.41) 78.88 ($\downarrow$0.77)
Corp.Issuers 68.31 ($\uparrow$14.24) 74.32 ($\uparrow$4.65) 79.26 ($\uparrow$16.08) 79.36 ($\uparrow$4.94) 87.02 ($\downarrow$1.16) 77.23 ($\uparrow$24.42) 75.58 ($\uparrow$9.88) 60.08 ($\downarrow$18.31) 79.07 ($\downarrow$0.58)
Table 1: Comparison of accuracy (%) of prompting techniques. ‘FinCoT’ simultaneously applies expert reasoning blueprints from all CFA domains, while each ‘(DomainName)’ (e.g., ‘Economics’) row applies domain-specific blueprints individually. ($\uparrow$/$\downarrow$) Denote accuracy improvement or decline relative to the SP baseline, colored green for ($\uparrow$) and red for ($\downarrow$). Bold values highlight the best-performing prompt variant for each model. (*) Indicates that the accuracy improvement among the model-level prompt variants is statistically significant ($p<0.05$) based on paired bootstrap testing; domain-specific rows are not tested for significance.
@@
Table˜1 reports zero-shot accuracies for four prompting strategies (SP, UST-CoT, ST-CoT, FinCoT) across our model suite. Under the basic SP prompt, the instruction-tuned Qwen3-8B (Thinker) attains the highest accuracy among general-purpose models (88.18%), while the financial model Fin-o1 (8B) leads its group at 79.65%. These strong baselines clearly highlight the effectiveness of instruction tuning and domain specificity.
Pretrained Models:
On Qwen2.5-7B, FinCoT (All Blueprints) yields a +7.95 pp improvement over SP (95% CI [6.30, 9.59], $p<0.001$), while UST-CoT and ST-CoT perform worse than the baseline. When FinCoT is applied using a single-domain blueprint (e.g., Financial Reporting), the gains increase substantially to +15.21 pp. Similarly, on Qwen3-8B-Base, FinCoT delivers the strongest overall boost (+17.35 pp, 95% CI [15.02, 19.77], $p<0.001$). These findings further underscore the importance of structured domain knowledge, particularly for models lacking instruction or domain alignment.
Instruction Models:
We evaluate three prominent instruction-tuned variants. On Qwen2.5-7B-Instruct, FinCoT improves accuracy by +4.55 pp over SP, compared to -1.46 pp with UST-CoT and -0.3 pp with ST-CoT. On Qwen3-8B (Thinker), FinCoT yields a slight drop (–0.96 pp), while ST-CoT shows no change and UST-CoT yields a modest +0.87 pp. Gemma-3-12B-IT, a strong instruction-tuned baseline (52.81% SP), benefits substantially from all strategies: +25.00 pp (UST-CoT), +23.93 pp (ST-CoT), and +22.77 pp (FinCoT). Notably, domain-specific FinCoT prompts (e.g., Derivatives, Corporate Issuers) provide even larger boosts (+24.42 pp), indicating that blueprint reasoning complements instruction tuning by addressing specialized financial gaps.
Financial-Specific Models:
FinCoT also helps specialized models like Fin-R1, confirming that blueprint prompting provides complementary gains beyond fine-tuning. However, for models with strong built-in reasoning, such as DianJin-R1-7B and Fin-o1-8B, FinCoT offers limited improvement or slight degradation—likely due to conflicts between external scaffolds and internal reasoning routines. These outcomes suggest diminishing returns for CoT prompting when domain alignment and reasoning are already deeply encoded.
Overall, FinCoT is most impactful for models lacking prior task-specific adaptation. By grounding reasoning in structured financial workflows, it bridges key gaps in zero-shot settings without requiring additional tuning. This pattern highlights a trade-off between model internalization and prompt-time controllability. Future work could explore hybrid strategies that adapt prompting depth based on model alignment.
Figure 4: Accuracy improvements (%) of each FinCoT domain-specific prompt compared to Standard Prompting (SP). Subfigure (a) shows results on Qwen3-8B-Base (pretrained), while (b) shows results on Fin-o1-8B (finance-specific).